# Less is More: One-shot-subgraph Link Prediction on Large-scale Knowledge Graphs

### v1

@@ 4.3 Theory: The extrapolation power of one-shot-subgraph link prediction
Further, we investigate the extrapolation power of link prediction across graph scales, i.e., training and inference on different scales of graphs. For example, training on small subgraphs $\mathcal{G}^{\text{train}}_{s}$ and testing on large subgraphs $\mathcal{G}^{\text{test}}_{s}$, where the ratio of entities $\nicefrac{{|\mathcal{V}^{\text{test}}_{s}|}}{{|\mathcal{V}^{\text{train}}_{s}|}}\!\gg\!1$. This scenario is practical for recommendation systems on social networks that can encounter much larger graphs in the testing phase. Intuitively, training on smaller $\mathcal{G}^{\text{train}}_{s}$ can save time for its faster convergence on subgraphs (Shi et al., 2023), while predicting on larger $\mathcal{G}^{\text{test}}_{s}$ might gain promotion for more support of facts in $\mathcal{G}^{\text{test}}_{s}$. Nonetheless, the Theorem 1 below proves that link prediction can become unreliable as the test graph grows. That is, if we use a small subgraph for training and a large subgraph for testing, the predictor will struggle to give different predictions within and across sampling distributions by $g$, even when these probabilities are arbitrarily different in the underlying graph model. Our empirical results in Fig. 4 support this theoretical finding. Hence, it is necessary to strike a balance of subgraphs’ scale.
@@
[…]
@@ 5 Experiments
Metrics. We adopt the filtered ranking-based metrics for evaluation, i.e., mean reciprocal ranking (MRR) and Hit@$k$ (i.e., both Hit@$1$ and Hit@$10$), following (Bordes et al., 2013; Teru et al., 2020; Wang et al., 2017; Zhu et al., 2021). For both metrics, a higher value indicates a better performance. Main Results. As results shown in Tab. 1 and Tab. 2, our one-shot-subgraph link prediction method achieves leading performances on all five large-scale benchmarks over all the baselines. Especially on the largest OGBL-WIKIKG2 dataset, a $16.6\%$ improvement in Test MRR can be achieved. Note the results attribute to a deep GNN (high expressiveness) and small subgraphs (essential information) extracted by sampling $10\%$ of entities on average for answering specific queries. Which means, it is unnecessary to utilize the whole KG in link prediction; meanwhile, only a small proportion of entities and facts are essential for answering specific queries that can be quickly identified by the PPR heuristics. In what follows, we conduct an in-depth analysis of the properties of the proposed method. The Sampling Distribution. We empirically evaluate to what extent the entities relevant to a specific query can be identified by heuristics, e.g., BFS, RW, and PPR. We quantify their power of identifying potential answers via the metric of Coverage Ratio $\texttt{CR}\!=\!\nicefrac{{1}}{{|\mathcal{E}^{\text{test}}|}}\sum\nolimits_{(u,q,v)\in\mathcal{E}^{\text{test}}}\mathbb{I}\{v\!\in\!\mathcal{V}_{s}\}$, i.e., the ratio of covered answer entities that remain in the set of sampled entities $\mathcal{V}_{s}$. As shown in Tab. 3 and Fig. 3, PPR gets a much higher CR and notably outperforms other heuristics in identifying potential answers.
@@
Training and Inference Efficiency. Next, we conduct an efficiency study to investigate the improvement of efficiency brought by the proposed one-shot-subgraph link prediction framework. The running time and GPU memory of an $8$-layer GNN are summarized in Tab. 7. As can be seen, a notable advantage of decoupling is that it has less computing cost in both terms of less running time and also less memory cost. Particularly, on the YAGO3-10 dataset, the existing GNN-based methods will run out of memory with a deep architecture. However, with the subgraph sampling of lower ratios of $r^{q}_{\mathcal{V}}$ and $r^{q}_{\mathcal{E}}$, the learning and prediction of GNNs become feasible that is with less memory cost and achieving state-of-the-art performance. Hence, we show that our method is effective and also efficient that it supports the learning and prediction of deep GNNs on large-scale knowledge graphs. Besides, we provide a detailed efficiency comparison between our method (with 10% entities) and the original implementation (with 100% entities) on two SOTA methods, NBFNet and RED-GNN. Tab. 6 shows that the training time is significantly reduced when learning with our method. Notably, on the YAGO3-10 dataset, $94.3\%$ and $94.5\%$ of training time (for one epoch) can be saved for NBFNet and RED-GNN, respectively. Besides, our method boosts the performance as advantage 2, where the performance improvement can come from a deeper GNN and a smaller observation graph (detailed analysis in Appendix. D.3). Full evaluations and more discussions are elaborated in the Appendix. C.
@@ A.2 Complexity Analysis
Next, we compare three different manners from perspectives of computation and parameter. In a nutshell, semantic models are computation-efficient but parameter-expensive, while structural models are parameter-efficient but computation-expensive. Here, we aim to make the best of the both worlds by designing a framework that is parameter-efficient and computation-efficient. The proposed one-shot-subgraph link prediction models, as will be elaborated in Sec. 3, is with the new prediction manner that $\mathcal{G}\xmapsto{g_{\bm{\phi}},(u_{q},r_{q})}\mathcal{G}_{s}\xmapsto{f_{\bm{\theta}}}\hat{\bm{Y}}$. The key difference here is that, instead of the original $\mathcal{G}$, the predictor $f_{\bm{\theta}}$ is acting on the query-dependent subgraph $\mathcal{G}_{s}$ extracted by the sampler $g_{\bm{\phi}}$. By contrast, we show that the proposed subgraph models make the best of both worlds: (1) subgraph models are computation-efficient, as the extracted subgraph is much smaller than the original graph, i.e., $|\mathcal{V}_{s}|\!\ll\!|\mathcal{V}|$ and $|\mathcal{E}_{s}|\!\ll\!|\mathcal{E}|$; (2) inherits from Sadeghian et al. (2019); Zhu et al. (2021); Zhang & Yao (2022), subgraph models are also parameter-efficient: only requires the relations’ embeddings but not the expensive entities’ embeddings. By contrast, the semantic models need to learn entities’s embeddings, which are parameter-expensive and only applicable in transductive settings. A detailed comparison of parameter and computation complexity is summarized in Tab. 9.
@@ A.3 Proof for Theorem 1
Here, let $\mathcal{G}^{\text{train}}_{s}\!\sim\!\mathbb{P}_{\mathcal{G}}$ and $\mathcal{G}^{\text{test}}_{s}\!\sim\!\mathbb{P}_{\mathcal{G}}$ be the training and testing graphs that are sampled distribution $\mathbb{P}_{\mathcal{G}}$. Let $\mathcal{G}^{\text{test}}$ be large enough to satisfy $\nicefrac{{\sqrt{|\mathcal{V}_{s}^{\text{test}}|}}}{{\sqrt{\log(2|\mathcal{V}_{s}^{\text{test}}|/p)}}}\geq\nicefrac{{4\sqrt{2}}}{{d_{\min}}}$, where $d_{\min}$ is the constant of graphon degree (Diaconis & Janson, 2007). In this part, we follow Maskey et al. (2022); Zhou et al. (2022) with the definitions of graph message-passing neural network $\Phi$ and continuous message passing neural network $\Psi$. Note that $\Phi$ is equivalent to the message passing function of the predictor $f_{\bm{w}}$ that generates the entity representation $\bm{h}_{v}^{(\ell)}$ in each layer $\ell=1,2,\cdots,L$. Besides, $\Psi$ is the continuous counterpart of $\Phi$. It uses the continuous aggregation instead of the discrete one w.r.t. the discrete graph topology in $\Phi$. With the two above functions, $\lVert\Phi\lVert_{\infty},\lVert\Psi\lVert_{\infty}$ can be trained and determined by the original graph $\mathcal{G}^{\text{train}}$ and hyperparameters $r^{q}_{\mathcal{V}},r^{q}_{\mathcal{E}}$ and $L$. Following Maskey et al. (2022); Zhou et al. (2022), let $p\in(0,\nicefrac{{1}}{{\sum_{\ell=1}^{L}2|\bm{h}^{(\ell)}|+1}})$ we with probability at least $1\!-\!\sum_{\ell=1}^{L}2(|\bm{h}^{(\ell)}|\!+\!1)p$, the difference $\delta$ between $\lVert\Phi\lVert_{\infty}$ and $\lVert\Psi\lVert_{\infty}$ is bounded by
@@
The first inequality holds by the triangle inequality. In contrast, the second inequality holds since $\Psi_{u}\!=\!\Psi_{v}$. Note that $u$ and $v$ are isomorphic in the topology of the test graph (Zhou et al., 2022) and thus with the same representations. Next, with probability at least $1\!-\!\sum_{\ell=1}^{L}2(|\bm{h}^{(\ell)}|\!+\!1)p$, for arbitrary entity $v^{\prime}\in\mathcal{V}_{s}^{\text{test}}$, we have
@@ Appendix B Implementation Details
method
$\bm{m}_{(e_{s},r,e_{o})}^{(\ell)}:=\text{MESS}(\cdot)$
$\bm{h}_{e_{o}}^{(\ell)}:=\text{AGG}(\cdot)$
R-GCN (Schlichtkrull et al., 2018)
$\bm{W}^{(\ell)}_{r}\bm{h}^{(\ell-1)}_{e_{s}}$, where $\bm{W}_{r}$ depends on the relation $r$
$\bm{W}^{(\ell)}_{o}\bm{h}^{(\ell-1)}_{e_{o}}\!+\!\sum_{e_{o}\in\mathcal{N}(e_{s})}\frac{1}{c_{o,r}}\bm{m}_{(e_{s},r,e_{o})}^{(\ell)}$
CompGCN (Vashishth et al., 2019)
$\bm{W}^{(\ell)}_{\lambda(r)}\{-,*,\star\}(\bm{h}^{(\ell-1)}_{e_{s}},\bm{h}^{(\ell)}_{r})$, where $\bm{W}^{(\ell)}_{\lambda(r)}$ depends on the direction of $r$
$\sum_{e_{o}\in\mathcal{N}(e_{s})}\bm{m}_{(e_{s},r,e_{o})}^{(\ell)}$
GraIL (Teru et al., 2020)
${\alpha}^{(\ell)}_{(e_{s},r,e_{o})|r_{q}}(\bm{W}^{(\ell)}_{1}\bm{h}^{(\ell-1)}_{e_{s}}+\bm{W}^{(\ell)}_{2}\bm{h}_{e_{o}}^{(\ell-1)})$, where ${\alpha}^{(\ell)}_{(e_{s},r,e_{o})|r_{q}}$ is the attention weight.
$\bm{W}^{(\ell)}_{o}\bm{h}^{(\ell-1)}_{e_{o}}\!+\!\sum_{e_{o}\in\mathcal{N}(e_{s})}\frac{1}{c_{o,r}}\bm{m}_{(e_{s},r,e_{o})}^{(\ell)}$
NBFNet (Zhu et al., 2021)
$\bm{W}^{(\ell)}\{+,*,\circ\}(\bm{h}_{e_{s}}^{(\ell-1)},\bm{w}_{q}(r,r_{q}))$, where $\bm{w}_{q}(r,r_{q})$ is a query-dependent weight vector
@@
RED-GNN (Zhang & Yao, 2022)
$\alpha_{(e_{s},r,e_{o})|r_{q}}^{(\ell)}(\bm{h}^{(\ell-1)}_{e_{s}}+\bm{h}^{(\ell)}_{r})$, where $\alpha_{(e_{s},r,e_{o})|r_{q}}^{(\ell)}$ is the attention weight
$\sum_{e_{o}\in\mathcal{N}(e_{s})}\bm{m}_{(e_{s},r,e_{o})}^{(\ell)}$
Design space of the predictor. From the model’s perspective, we build the configuration space of the predictor and further utilize the advantages of existing structural models introduced in Sec. 2. Three query-dependent message functions MESS(normal-$\cdot$) are considered here, including DRUM (Sadeghian et al., 2019) (denoted as $\texttt{M}_{\text{DRUM}}$), NBFNet (Zhu et al., 2021) ($\texttt{M}_{\text{NBFNet}}$), and RED-GNN (Zhang & Yao, 2022) ($\texttt{M}_{\text{REDGNN}}$). The effective message is propagated from $u_{q}$ to the entities in subgraph $\mathcal{G}_{s}$. Generally, the message propagation can be formulated as
@@
DROPOUT()normal-⋅($\cdot$)( ⋅ ) ACT()normal-⋅($\cdot$)( ⋅ ) AGG()normal-⋅($\cdot$)( ⋅ ) MESS()normal-⋅($\cdot$)( ⋅ ) Dimension
@@
No. layers ($L$) Repre. initialization Layer-wise shortcut Repre. concatenation READOUT()normal-⋅($\cdot$)( ⋅ )

### v2

@@ 4.3 Theory: The extrapolation power of one-shot-subgraph link prediction
Further, we investigate the extrapolation power of link prediction across graph scales, i.e., training and inference on different scales of graphs. For example, training on small subgraphs $\mathcal{G}^{\text{train}}_{s}$ and testing on large subgraphs $\mathcal{G}^{\text{test}}_{s}$, where the ratio of entities $\nicefrac{{|\mathcal{V}^{\text{test}}_{s}|}}{{|\mathcal{V}^{\text{train}}_{s}|}}\!\gg\!1$. This scenario is practical for recommendation systems on social networks that can encounter much larger graphs in the testing phase. Intuitively, training on smaller $\mathcal{G}^{\text{train}}_{s}$ can save time for its faster convergence on subgraphs (Shi et al., 2023), while predicting on larger $\mathcal{G}^{\text{test}}_{s}$ might gain promotion for more support of facts in $\mathcal{G}^{\text{test}}_{s}$.
Nonetheless, the Theorem 1 below proves that link prediction can become unreliable as the test graph grows. That is, if we use a small subgraph for training and a large subgraph for testing, the predictor will struggle to give different predictions within and across sampling distributions by $g$, even when these probabilities are arbitrarily different in the underlying graph model. Our empirical results in Fig. 4 support this theoretical finding. Hence, it is necessary to strike a balance of subgraphs’ scale.
@@
Let $\mathcal{G}^{\text{train}}_{s}\!\sim\!\mathbb{P}_{\mathcal{G}}$ and $\mathcal{G}^{\text{test}}_{s}\!\sim\!\mathbb{P}_{\mathcal{G}}$ be the training and testing graphs that are sampled from distribution $\mathbb{P}_{\mathcal{G}}$. Consider any two test entities $u,v\!\in\!\mathcal{V}_{s}^{\text{test}}$, for which we can make a prediction decision of fact $(u,q,v)$ with the predictor $f_{\bm{\theta}}$, i.e., $\hat{\bm{y}}_{v}\!=\!f_{\bm{\theta}}(\mathcal{G}^{\text{test}}_{s})_{v}\!\neq\!\tau$. Let $\mathcal{G}^{\text{test}}$ be large enough to satisfy $\nicefrac{{\sqrt{|\mathcal{V}_{s}^{\text{test}}|}}}{{\sqrt{\log(2|\mathcal{V}_{s}^{\text{test}}|/p)}}}\geq\nicefrac{{4\sqrt{2}}}{{d_{\min}}}$, where $d_{\min}$ is the constant of graphon degree (Diaconis & Janson, 2007). Then, for an arbitrary threshold $\tau\!\in\![0,1]$, the testing subgraph $\mathcal{G}_{s}^{\text{test}}$ satisfies that
$\frac{\sqrt{|\mathcal{V}_{s}^{\text{test}}|}}{\sqrt{\log(2|\mathcal{V}_{s}^{\text{test}}|/p)}}\;\geq\;\frac{2(C_{1}+C_{2}\lVert g\lVert_{\infty})}{|f_{\bm{\theta}}(\mathcal{G}^{\text{test}}_{s})_{v}-\tau|/L(M^{\text{train}})}.$ (6)
where the underlying generative function of graph signal $g\!\in\!L^{\infty}$ is with the essential supreme norm as in (Maskey et al., 2022; Zhou et al., 2022). The $p,C_{1},C_{2}$ are constants and depend on $M^{\text{train}}$ where $\min(supp(|\mathcal{V}_{s}^{\text{test}}|))\!\gg\!M^{\text{train}}\!=\!\max(supp(|\mathcal{V}_{s}^{\text{train}}|))$. It means any test graph can be much larger than the largest possible training graph, and $supp$ indicates the support of a distribution. Then, if $u$ and $v$ are isomorphic in topology and with the same representations, we have a probability at least $1\!-\!\sum_{\ell=1}^{L}2(|\bm{h}^{(\ell)}|\!+\!1)p$ with hidden size $|\bm{h}^{(\ell)}|$ that the same predictions can be obtained whether $u,v$ are generated by the same or distinct $g$. The detailed proof can be found in Appendix. A.
@@ 5 Experiments
Metrics. We adopt the filtered ranking-based metrics for evaluation, i.e., mean reciprocal ranking (MRR) and Hit@$k$ (i.e., both Hit@$1$ and Hit@$10$), following (Bordes et al., 2013; Teru et al., 2020; Wang et al., 2017; Zhu et al., 2021). For both metrics, a higher value indicates a better performance.
Main Results. As results shown in Tab. 1 and Tab. 2, our one-shot-subgraph link prediction method achieves leading performances on all five large-scale benchmarks over all the baselines. Especially on the largest OGBL-WIKIKG2 dataset, a $16.6\%$ improvement in Test MRR can be achieved. Note the results attribute to a deep GNN (high expressiveness) and small subgraphs (essential information) extracted by sampling $10\%$ of entities on average for answering specific queries. Which means, it is unnecessary to utilize the whole KG in link prediction; meanwhile, only a small proportion of entities and facts are essential for answering specific queries that can be quickly identified by the PPR heuristics. In what follows, we conduct an in-depth analysis of the properties of the proposed method.
The Sampling Distribution. We empirically evaluate to what extent the entities relevant to a specific query can be identified by heuristics, e.g., BFS, RW, and PPR. We quantify their power of identifying potential answers via the metric of Coverage Ratio $\texttt{CR}\!=\!\nicefrac{{1}}{{|\mathcal{E}^{\text{test}}|}}\sum\nolimits_{(u,q,v)\in\mathcal{E}^{\text{test}}}\mathbb{I}\{v\!\in\!\mathcal{V}_{s}\}$, i.e., the ratio of covered answer entities that remain in the set of sampled entities $\mathcal{V}_{s}$. As shown in Tab. 3 and Fig. 3, PPR gets a much higher CR and notably outperforms other heuristics in identifying potential answers.
@@
Training and Inference Efficiency. Next, we conduct an efficiency study to investigate the improvement of efficiency brought by the proposed one-shot-subgraph link prediction framework. The running time and GPU memory of an $8$-layer GNN are summarized in Tab. 7. As can be seen, a notable advantage of decoupling is that it has less computing cost in both terms of less running time and also less memory cost. Particularly, on the YAGO3-10 dataset, the existing GNN-based methods will run out of memory with a deep architecture. However, with the subgraph sampling of lower ratios of $r^{q}_{\mathcal{V}}$ and $r^{q}_{\mathcal{E}}$, the learning and prediction of GNNs become feasible that is with less memory cost and achieving state-of-the-art performance. Hence, we show that our method is effective and also efficient that it supports the learning and prediction of deep GNNs on large-scale knowledge graphs.
Besides, we provide a detailed efficiency comparison between our method (with 10% entities) and the original implementation (with 100% entities) on two SOTA methods, NBFNet and RED-GNN. Tab. 6 shows that the training time is significantly reduced when learning with our method. Notably, on the YAGO3-10 dataset, $94.3\%$ and $94.5\%$ of training time (for one epoch) can be saved for NBFNet and RED-GNN, respectively. Besides, our method boosts the performance as advantage 2, where the performance improvement can come from a deeper GNN and a smaller observation graph (detailed analysis in Appendix. D.3). Full evaluations and more discussions are elaborated in the Appendix. C.
@@ A.2 Complexity Analysis
Next, we compare three different manners from perspectives of computation and parameter.
In a nutshell, semantic models are computation-efficient but parameter-expensive, while structural models are parameter-efficient but computation-expensive. Here, we aim to make the best of the both worlds by designing a framework that is parameter-efficient and computation-efficient. The proposed one-shot-subgraph link prediction models, as will be elaborated in Sec. 3, is with the new prediction manner that $\mathcal{G}\xmapsto{g_{\bm{\phi}},(u_{q},r_{q})}\mathcal{G}_{s}\xmapsto{f_{\bm{\theta}}}\hat{\bm{Y}}$. The key difference here is that, instead of the original $\mathcal{G}$, the predictor $f_{\bm{\theta}}$ is acting on the query-dependent subgraph $\mathcal{G}_{s}$ extracted by the sampler $g_{\bm{\phi}}$.
By contrast, we show that the proposed subgraph models make the best of both worlds: (1) subgraph models are computation-efficient, as the extracted subgraph is much smaller than the original graph, i.e., $|\mathcal{V}_{s}|\!\ll\!|\mathcal{V}|$ and $|\mathcal{E}_{s}|\!\ll\!|\mathcal{E}|$; (2) inherits from Sadeghian et al. (2019); Zhu et al. (2021); Zhang & Yao (2022), subgraph models are also parameter-efficient: only requires the relations’ embeddings but not the expensive entities’ embeddings. By contrast, the semantic models need to learn entities’s embeddings, which are parameter-expensive and only applicable in transductive settings.
A detailed comparison of parameter and computation complexity is summarized in Tab. 9.
@@ A.3 Proof for Theorem 1
Here, let $\mathcal{G}^{\text{train}}_{s}\!\sim\!\mathbb{P}_{\mathcal{G}}$ and $\mathcal{G}^{\text{test}}_{s}\!\sim\!\mathbb{P}_{\mathcal{G}}$ be the training and testing graphs that are sampled distribution $\mathbb{P}_{\mathcal{G}}$. Let $\mathcal{G}^{\text{test}}$ be large enough to satisfy $\nicefrac{{\sqrt{|\mathcal{V}_{s}^{\text{test}}|}}}{{\sqrt{\log(2|\mathcal{V}_{s}^{\text{test}}|/p)}}}\geq\nicefrac{{4\sqrt{2}}}{{d_{\min}}}$, where $d_{\min}$ is the constant of graphon degree (Diaconis & Janson, 2007).
In this part, we follow Maskey et al. (2022); Zhou et al. (2022) with the definitions of graph message-passing neural network $\Phi$ and continuous message passing neural network $\Psi$. Note that $\Phi$ is equivalent to the message passing function of the predictor $f_{\bm{w}}$ that generates the entity representation $\bm{h}_{v}^{(\ell)}$ in each layer $\ell=1,2,\cdots,L$. Besides, $\Psi$ is the continuous counterpart of $\Phi$. It uses the continuous aggregation instead of the discrete one w.r.t. the discrete graph topology in $\Phi$.
With the two above functions, $\lVert\Phi\lVert_{\infty},\lVert\Psi\lVert_{\infty}$ can be trained and determined by the original graph $\mathcal{G}^{\text{train}}$ and hyperparameters $r^{q}_{\mathcal{V}},r^{q}_{\mathcal{E}}$ and $L$. Following Maskey et al. (2022); Zhou et al. (2022), let $p\in(0,\nicefrac{{1}}{{\sum_{\ell=1}^{L}2|\bm{h}^{(\ell)}|+1}})$ we with probability at least $1\!-\!\sum_{\ell=1}^{L}2(|\bm{h}^{(\ell)}|\!+\!1)p$, the difference $\delta$ between $\lVert\Phi\lVert_{\infty}$ and $\lVert\Psi\lVert_{\infty}$ is bounded by
@@
The first inequality holds by the triangle inequality. In contrast, the second inequality holds since $\Psi_{u}\!=\!\Psi_{v}$. Note that $u$ and $v$ are isomorphic in the topology of the test graph (Zhou et al., 2022) and thus with the same representations.
Next, with probability at least $1\!-\!\sum_{\ell=1}^{L}2(|\bm{h}^{(\ell)}|\!+\!1)p$, for arbitrary entity $v^{\prime}\in\mathcal{V}_{s}^{\text{test}}$, we have
@@ Appendix B Implementation Details
method $\bm{m}_{(e_{s},r,e_{o})}^{(\ell)}:=\text{MESS}(\cdot)$ $\bm{h}_{e_{o}}^{(\ell)}:=\text{AGG}(\cdot)$
R-GCN (Schlichtkrull et al., 2018) $\bm{W}^{(\ell)}_{r}\bm{h}^{(\ell-1)}_{e_{s}}$, where $\bm{W}_{r}$ depends on the relation $r$ $\bm{W}^{(\ell)}_{o}\bm{h}^{(\ell-1)}_{e_{o}}\!+\!\sum_{e_{o}\in\mathcal{N}(e_{s})}\frac{1}{c_{o,r}}\bm{m}_{(e_{s},r,e_{o})}^{(\ell)}$
CompGCN (Vashishth et al., 2019) $\bm{W}^{(\ell)}_{\lambda(r)}\{-,*,\star\}(\bm{h}^{(\ell-1)}_{e_{s}},\bm{h}^{(\ell)}_{r})$, where $\bm{W}^{(\ell)}_{\lambda(r)}$ depends on the direction of $r$ $\sum_{e_{o}\in\mathcal{N}(e_{s})}\bm{m}_{(e_{s},r,e_{o})}^{(\ell)}$
GraIL (Teru et al., 2020) ${\alpha}^{(\ell)}_{(e_{s},r,e_{o})|r_{q}}(\bm{W}^{(\ell)}_{1}\bm{h}^{(\ell-1)}_{e_{s}}+\bm{W}^{(\ell)}_{2}\bm{h}_{e_{o}}^{(\ell-1)})$, where ${\alpha}^{(\ell)}_{(e_{s},r,e_{o})|r_{q}}$ is the attention weight. $\bm{W}^{(\ell)}_{o}\bm{h}^{(\ell-1)}_{e_{o}}\!+\!\sum_{e_{o}\in\mathcal{N}(e_{s})}\frac{1}{c_{o,r}}\bm{m}_{(e_{s},r,e_{o})}^{(\ell)}$
@@
RED-GNN (Zhang & Yao, 2022) $\alpha_{(e_{s},r,e_{o})|r_{q}}^{(\ell)}(\bm{h}^{(\ell-1)}_{e_{s}}+\bm{h}^{(\ell)}_{r})$, where $\alpha_{(e_{s},r,e_{o})|r_{q}}^{(\ell)}$ is the attention weight $\sum_{e_{o}\in\mathcal{N}(e_{s})}\bm{m}_{(e_{s},r,e_{o})}^{(\ell)}$
Design space of the predictor. From the model’s perspective, we build the configuration space of the predictor and further utilize the advantages of existing structural models introduced in Sec. 2. Three query-dependent message functions MESS($\cdot$) are considered here, including DRUM (Sadeghian et al., 2019) (denoted as $\texttt{M}_{\text{DRUM}}$), NBFNet (Zhu et al., 2021) ($\texttt{M}_{\text{NBFNet}}$), and RED-GNN (Zhang & Yao, 2022) ($\texttt{M}_{\text{REDGNN}}$). The effective message is propagated from $u_{q}$ to the entities in subgraph $\mathcal{G}_{s}$.
Generally, the message propagation can be formulated as
@@
DROPOUT()($\cdot$)( ⋅ ) ACT()($\cdot$)( ⋅ ) AGG()($\cdot$)( ⋅ ) MESS()($\cdot$)( ⋅ ) Dimension
@@
No. layers ($L$) Repre. initialization Layer-wise shortcut Repre. concatenation READOUT()($\cdot$)( ⋅ )